A company is expected to have earnings per share of $3.99 this year and to pay a dividend of $2.14. The discount rate for the stock is 13.7% and the rate of return on reinvested earnings is 22.9%. What is the sustainable growth rate? Enter your answer as a percentage. Do not include the percentage sign in your answer.

Enter your response below rounded to 2 DECIMAL PLACES.
Answer #1

Answer:

Payout Ratio = Dividend Per share / Earning per share
Payout Ratio = $2.14 / $3.99
Payout Ratio = 0.5363

Retention Ratio = 1- Payout Ratio
Retention Ratio = 1 – 0.5363
Retention Ratio = 0.4637

Sustainable Growth Rate = (ROE * b) / (1 – ROE *b)
Sustainable Growth Rate = (0.229 * 0.4637) / (1 – 0.229 * 0.4637)
Sustainable Growth Rate = 0.1062 / (1 – 0.1062)
Sustainable Growth Rate = 0.1062 / 0.8938
Sustainable Growth Rate = 0.1188 or 11.88%